# Rate-Adaptive Generative Semantic Communication Using Conditional Diffusion Models

### v1

@@ I Introduction
Recently, the integration of deep learning into wireless communication system designs has gained traction, driven by the exceptional information processing capabilities of various deep learning models. In this context, deep learning-based JSCC (DJSCC) has stimulated significant interest, particularly through the use of autoencoders (AEs) and their variants, such as variational AEs (VAEs). DJSCC maps input images into low-dimensional symbol vectors for transmission. A pioneering work in this area is DeepJSCC[2], which is built on an AE-based framework that allows for joint optimization and mitigates the cliff-edge effect commonly observed in traditional separation-based schemes. Moreover, the authors in [4] introduced a Transformer-based framework that enhances image fidelity by incorporating channel feedback to adaptively reconstruct images under varying wireless conditions. Inspired by entropy model-based source coding [5], the authors in [6] proposed a rate-adaptive JSCC system, where the number of transmitted symbols is determined by the entropy estimated by the entropy models. This work was further evolved in [7] into a digital version, enabling practical application in modern digital systems.
@@
Despite the superior performance of current DJSCC transmission systems, they are typically optimized for mean-square error (MSE) distortion metrics such as peak signal-to-noise ratio (PSNR), which assesses pixel-level similarity between reconstructed and source images. However, it is increasingly recognized that high pixel-level similarity does not necessarily indicate high perceptual quality, which reflects how humans perceive the image’s realism and visual appeal [8]. To improve the perceptual quality of transmitted images, the authors in [9] introduced adversarial and perceptual losses into DJSCC. Moreover, [10] developed a dual-path framework to extract both pixel-level information and perceptual information. Alternatively, [11] combined DJSCC with the denoising diffusion probabilistic model (DDPM) [12], a powerful generative model that creates images from Gaussian noise. In [11], the reconstructed image is split into two components: range-space and null-space. The range-space that captures the primary structure is transmitted via DJSCC, while the null-space, refining details, is generated at the receiver using diffusion models. Another approach, presented in [13], leveraged diffusion models by first performing an initial reconstruction with DJSCC, followed by a neural network to guide the denoising process. Although [11] and [13] have achieved performance improvements, the decoding process at the receiver is time-consuming, as denoising requires hundreds of steps. Additionally, the training is performed module by module without joint optimization, potentially leading to performance loss. Besides, they typically support fixed-rate transmission, lacking the ability to determine optimal bandwidth for each image.
@@
In this letter, we introduce a novel framework called conditional diffusion models-based generative DJSCC (CDM-JSCC) for wireless image transmission. Unlike previous methods that rely on DDPM noise prediction (referred to as $\epsilon$-prediction)[11, 13], our method utilizes $\mathcal{X}$-prediction [14] to directly predict the source image. $\mathcal{X}$-prediction offers comparable performance to $\epsilon$-prediction but with significantly fewer steps, as its objective function resembles an autoencoder loss, enabling data reconstruction in a single iteration. Moreover, we optimize transmission bandwidth by utilizing entropy models to estimate the entropy of transmitted symbols. At the receiver, the transmitted symbols are utilized as conditional information to guide the conditional diffusion decoder to reconstruct images. Recently, the mamba-like linear attention (MLLA) technique has demonstrated superior performance in image processing tasks, benefiting from parallel computation and fast inference [15]. Building on this, we design our JSCC encoder and decoder with MLLA as the core architecture. Furthermore, we propose a multi-stage training strategy to achieve joint optimization, resulting in substantial enhancements in perceptual quality.
@@ II-A Rate-Adaptive JSCC
In this way, we are able to control the channel bandwidth ratio (CBR) based on the estimated entropy of $\bm{\tilde{z}}$. Specifically, $\bm{z}$ comprises multiple embedding vectors $z_{i}$ of length $C$. The JSCC encoder $f_{e}$ adjusts the length of each vector to ${k}_{i}=Q\left(-\beta\log p_{\tilde{z}_{i}|\bm{\tilde{y}}}\left(\tilde{z}_{i}|\bm{\tilde{y}}\right)\right)$, where $Q$ denotes the quantization operation, $\beta$ controls the relation between the prior $p_{\hat{z}_{i}|\bm{\tilde{y}}}\left(\hat{z}_{i}|\bm{\tilde{y}}\right)$, and the expected length of the corresponding symbol vector $s_{i}$. In particular, when the entropy of $\tilde{z}_{i}$ is high, a higher CBR is adopted, resulting in a larger $k_{i}$. This process is expressed as $\bm{s}=f_{e}(\bm{z},p_{\tilde{\bm{z}}|\bm{\tilde{y}}}\left(\tilde{\bm{z}}|\bm{\tilde{y}}\right);\bm{\theta}_{f})\in\mathbb{C}^{k},$ where $\bm{s}$ denotes the channel input symbols, $k$ is the number of transmitted symbols, and $\bm{\theta}_{f}$ denotes the trainable parameters. Additionally, to meet the energy constraints of real-world communication systems, we ensure that $\bm{s}$ satisfies an average power constraint before transmission.
Then, $\bm{s}$ is transmitted through a noisy wireless channel, which is denoted by $\eta$. As the additive white Gaussian noise (AWGN) channel is adopted in our work, this process follows $\bm{\hat{s}}\triangleq\eta(\bm{s})=\bm{s}+\bm{n}$, where $\bm{n}\sim\mathcal{CN}(0,\sigma^{2}\bm{I}_{k\times k})$ is a complex Gaussian vector with variance $\sigma^{2}$. At the receiver, $\bm{\hat{s}}$ is first decoded to obtain the reconstructed latent $\bm{\hat{z}}=f_{d}(\bm{\hat{s}};\bm{\phi}_{f})$, where $f_{d}$ represents the JSCC decoder, and $\bm{\phi}_{f}$ denotes its trainable parameters. The receiver utilizes $\bm{\hat{z}}$ as an additional “content” latent to guide the conditional diffusion decoder $g_{d}$ to reconstruct images $\bm{\hat{x}}_{0}=g_{d}(\bm{\hat{x}}_{N},\bm{\hat{z}};\bm{\phi}_{g})$, where $\bm{\hat{x}}_{N}$ is the randomly sampled Gaussian source and $\bm{\phi}_{g}$ represents the trainable parameters, as shown in Fig. 1. The details of the conditional diffusion decoder will be provided in the following section.
@@ III-A Model Architecture
III-A1 Transmitter
The transmitter consists of an entropy encoder and a rate-adaptive JSCC encoder. The entropy encoder comprises a hyperprior model to effectively capture spatial dependencies in the latent representation [5]. The encoder $g_{e}$ consists of $4$ residual blocks followed by $4$ down-sampling convolution layers connected sequentially. The hyperprior encoder is composed of a down-sampling convolution layer followed by a ReLU activation function, while the hyperprior decoder is composed of an up-sampling convolution layer and a ReLU activation function. Then, $\bm{z}$ is input to the JSCC encoder, where it is compressed into a variable-length latent representation based on the estimated entropy $p(\bm{\tilde{z}}|\bm{\tilde{y}})$. Specifically, elements are selected and masked sequentially in a one-dimensional checkerboard pattern, which captures global features more effectively than random selection or linear sequencing. In addition, our proposed model is built on the advanced MLLA skeleton, known for its exceptional performance in image processing tasks and fast inference speed. The JSCC encoder consists of $3$ MLLA blocks and multiple multi-layer perceptrons (MLPs).
III-A2 Receiver
The receiver consists of a rate-adaptive JSCC decoder and a conditional diffusion decoder. The JSCC decoder inverts the operations performed by the JSCC encoder. Then, $\bm{\hat{z}}$ is input to the diffusion decoder as conditional information to guide the denoising process. The diffusion denoising unit is built on a U-Net architecture, featuring skip connections that allow the output from each down-sampling unit to be directly passed to the corresponding up-sampling unit. The architecture includes $6$ down-sampling units and $6$ up-sampling units, with a mid-block in between. Each down-sampling unit, up-sampling unit, and mid-block is composed of residual blocks and linear attention mechanisms. In addition, $\bm{\hat{z}}$ is encoded into features of different scales by $4$ layers of hierarchical encoders. These features are concatenated with the output of the preceding down-sampling unit and embed time information as input for the next down-sampling unit. During testing, the output $\bm{\hat{x}}_{n-1}$ is passed through the diffusion denoising unit iteratively until $\bm{\hat{x}}_{0}$ is produced.
@@ III-B Loss Function
In a typical end-to-end JSCC system, the optimization focuses sorely on the MSE between the source image and the reconstructed image, with the loss function defined as:
$\mathcal{L}_{D}=\mathbb{E}_{\bm{x}_{0}}||\bm{x}_{0}-\bm{\hat{x}}_{0}||^{2},$ (8)
which aligns with the training objective of our conditional diffusion models as defined in (8), with the exception that step $n$ and noise $\epsilon$ are randomly sampled.
Moreover, for an entropy model-based rate-adaptive JSCC system, the loss function is typically defined as:
@@ III-C Training Strategy
To ensure training stability and enhance overall performance, we propose a multi-stage training strategy. Initially, we train each module individually to reduce complexity and facilitate convergence. Once all modules have been trained separately, we fine-tune the entire model to optimize performance. The complete multi-stage training process is detailed in Algorithm 1.
@@
3 First Phase: Train Compression Modules.
@@ 7       Calculate the loss function:
8       $\mathcal{L}=(1-\eta)\mathbb{E}_{\bm{x}_{0}}||\bm{x}_{0}-\bm{\bar{x}}_{0}||^{2}+\eta\mathbb{E}_{\bm{x}_{0}}[d_{p}(\bm{x}_{0},\bm{\bar{x}}_{0})]+\lambda{\mathbb{E}_{\bm{x}_{0}}[-\log p(\bm{\tilde{z}}|\bm{\tilde{y}})-\log p(\bm{\tilde{y}})]},$
@@
13 Second Phase: Train Transmission Modules.
@@
22 Third Phase: Finetune the whole model.
23 Load the parameters trained in the first and second stages.
@@ IV-A1 Basic Settings
We evaluate our CDM-JSCC on the Kodak image dataset, which consists of $24$ images, each with a resolution of $512\times 768\times 3$. During testing, we only employ $30$ denoising steps at the receiver, ensuring an efficient and fast decoding process. All experiments are conducted using Pytorch. For the training process, we use the Adam optimizer for stochastic gradient descent, starting with a learning rate of $1\times 10^{-4}$ and planning to reduce it after several epochs. The training dataset comprises $50,000$ randomly sampled images from the ImageNet dataset, which are randomly cropped to $256\times 256$. The batch size is set to $4$. We set $\lambda=0.0512$ and $\eta=0.5$ for the experiments.
Figure 3: The performance of the proposed schemes versus SNR.
@@ IV-A3 Considered Metrics
We evaluate the performance of our method along with benchmarks using typical metric PSNR as well as perceptual metrics LPIPS[16], and Fréchet Inception Distance (FID). PSNR calculates pixel-wise MSE distortion. LPIPS measures the $l$2 distance between two latent embeddings extracted by a pre-trained network [16]. FID calculates the distribution of source images and reconstructed images in the feature space using the Fréchet distance [17].
@@ IV-B Performance Comparison
We evaluate the performance of our proposed model on the AWGN channel, a representative and widely adopted method in this research community. In all subsequent experiments, the model is trained on a single SNR value and tested on the same SNR. In addition, by carefully adjusting the hyper-parameters, we control the CBR $=1/48$.
Fig. 3 compares CDM-JSCC with benchmarks across various SNRs, illustrating a comparative analysis of our proposed CDM-JSCC against the benchmarks. Notably, our model outperforms others across all perceptual quality metrics, including LPIPS and FID. This enhancement is attributed to the proposed loss function and the powerful generative diffusion models. However, it is important to note that our scheme shows some performance degradation in the PSNR metric compared to the MSE-optimized DeepJSCC and traditional separation-based schemes. This is because generative models tend to generate realistic and clear images and may overlook pixel-wise fidelity sometimes. Despite this, our model still outperforms GAN-JSCC in the PSNR metric. Unlike GAN-JSCC, our model does not plateau as SNR increases, maintaining the potential to transmit clear, realistic, and high-fidelity images at higher SNR levels.

### v3

@@ I Introduction
Recently, the integration of deep learning into wireless communication system designs has gained traction, driven by the exceptional information processing capabilities of various deep learning models. In this context, deep learning-based JSCC (DJSCC) has stimulated significant interest, particularly through the use of autoencoders (AEs) and their variants variational AEs. DJSCC maps input images into low-dimensional vectors for transmission. A pioneering work in this area is DeepJSCC[3], which is built on an AE-based framework that allows for joint optimization and mitigates the cliff-edge effect commonly observed in traditional separation-based schemes. Moreover, the authors in [4] introduced a Transformer-based framework that enhances image fidelity by incorporating channel feedback to adaptively reconstruct images under varying wireless conditions. Inspired by entropy model-based source coding [5], the authors in [6] proposed a rate-adaptive JSCC system, where the number of transmitted symbols is determined by the entropy estimated by the entropy models. This work was further evolved in [7] into a digital version, enabling practical application in modern digital systems.
@@
Despite the superior performance of current DJSCC transmission systems, they are typically optimized for mean-square error (MSE)-based distortion metrics such as peak signal-to-noise ratio (PSNR), which assesses pixel-level similarity between reconstructed and source images. However, it is increasingly recognized that high pixel-level similarity does not necessarily indicate high perceptual quality, which reflects how humans perceive the image’s realism and visual appeal [8]. Generative models, including generative adversarial networks (GANs) and denoising diffusion probabilistic models (DDPM) [9], excel at generating images with high perceptual quality from a predefined prior distribution, demonstrating their groundbreaking potential in generative semantic communications [10, 11, 12, 13, 14, 15, 16]. To improve the perceptual quality of transmitted images, the authors in [10] incorporated perceptual loss into a GAN-based DJSCC system. In [11], the authors leveraged DDPM to mitigate channel noise in wireless communications. Furthermore, in [14], DJSCC was employed for initial reconstruction, which was then refined by DDPM to further improve image quality. Another approach in [15] divided the reconstructed image into two components: range-space and null-space. The range-space that captures the primary structure is transmitted via DJSCC, while the null-space, refining details, is generated at the receiver using diffusion models. Although these methods achieved performance improvements, the decoding process at the receiver is time-consuming, as denoising requires hundreds of steps. Additionally, the training is performed module by module without joint optimization, potentially leading to performance loss. Besides, they typically support fixed-rate transmission, lacking the ability to determine optimal bandwidth for each image.
@@
In this letter, we introduce a novel framework called conditional diffusion models-based generative DJSCC (CDM-JSCC) for wireless image transmission. Unlike previous methods that rely on DDPM noise prediction (referred to as $\epsilon$-prediction)[15], our method utilizes $\mathcal{X}$-prediction [17] to directly predict the source image. $\mathcal{X}$-prediction offers comparable performance to $\epsilon$-prediction but with significantly fewer steps, as its objective function resembles an autoencoder loss, enabling data reconstruction in a single iteration. Moreover, we optimize transmission bandwidth by utilizing entropy models to estimate the entropy of transmitted symbols. At the receiver, the transmitted symbols are utilized as conditional information to guide the conditional diffusion decoder to reconstruct images. Recently, the mamba-like linear attention (MLLA) technique has demonstrated superior performance in image processing tasks, benefiting from parallel computation and fast inference [18]. Building on this, we design our JSCC encoder and decoder with MLLA as the core architecture. Furthermore, we propose a multi-stage training strategy to achieve joint optimization, resulting in substantial enhancements in perceptual quality.
@@ II-A Rate-Adaptive JSCC
In this way, we are able to control the channel bandwidth ratio (CBR) for each image based on the estimated entropy of $\bm{\tilde{z}}$. Images with rich details, characterized by higher entropy, may require more symbols for transmission. Specifically, $\bm{z}$ comprises multiple embedding vectors $z_{i}$ of length $C$. The JSCC encoder $f_{e}$ adjusts the length of each vector to ${k}_{i}=Q\left(-\beta\log p_{\tilde{z}_{i}|\bm{\tilde{y}}}\left(\tilde{z}_{i}|\bm{\tilde{y}}\right)\right)$, where $Q$ denotes the quantization operation, $\beta$ controls the relation between the prior $p_{\hat{z}_{i}|\bm{\tilde{y}}}\left(\hat{z}_{i}|\bm{\tilde{y}}\right)$, and the expected length of the corresponding symbol vector $s_{i}$. In particular, when the entropy of $\tilde{z}_{i}$ is high, a higher CBR is adopted, resulting in a larger $k_{i}$. This process is expressed as $\bm{s}=f_{e}(\bm{z},p_{\tilde{\bm{z}}|\bm{\tilde{y}}}\left(\tilde{\bm{z}}|\bm{\tilde{y}}\right);\bm{\theta}_{f})\in\mathbb{C}^{k},$ where $\bm{s}$ denotes the channel input symbols, $k$ is the number of transmitted symbols, and $\bm{\theta}_{f}$ denotes the trainable parameters. Additionally, to meet the energy constraints of real-world communication systems, we ensure that $\bm{s}$ satisfies an average power constraint before transmission.
Then, $\bm{s}$ is transmitted through a noisy wireless channel, which is denoted by $\eta$. As the additive white Gaussian noise (AWGN) channel is adopted in our work, this process follows $\bm{\hat{s}}\triangleq\eta(\bm{s})=\bm{s}+\bm{n}$, where $\bm{n}\sim\mathcal{CN}(0,\sigma^{2}\bm{I}_{k\times k})$ is a complex Gaussian vector with variance $\sigma^{2}$. At the receiver, $\bm{\hat{s}}$ is first decoded to obtain the reconstructed latent $\bm{\hat{z}}=f_{d}(\bm{\hat{s}};\bm{\phi}_{f})$, where $f_{d}$ represents the JSCC decoder, and $\bm{\phi}_{f}$ denotes its trainable parameters. The receiver utilizes $\bm{\hat{z}}$ as an additional “content” latent to guide the conditional diffusion decoder $g_{d}$ to reconstruct image $\bm{\hat{x}}_{0}=g_{d}(\bm{\hat{x}}_{N},\bm{\hat{z}};\bm{\phi}_{g})$, where $\bm{\hat{x}}_{N}$ is the randomly sampled Gaussian source and $\bm{\phi}_{g}$ represents the trainable parameters, as shown in Fig. 1. The details of the conditional diffusion decoder will be provided in the following section.
@@ III-A Model Architecture
The transmitter comprises an entropy encoder and a rate-adaptive JSCC encoder. The entropy encoder incorporates a hyperprior model to capture spatial dependencies in the latent representation [5]. Meanwhile, the JSCC encoder, built on the MLLA skeleton, compresses the fixed-length latent representation $\bm{z}$ into a variable-length output $\bm{s}$ using the estimated entropy $p(\bm{\tilde{z}}|\bm{\tilde{y}})$. Elements are selected sequentially in a one-dimensional checkerboard pattern to efficiently capture global features.
The receiver comprises a rate-adaptive JSCC decoder and a conditional diffusion decoder. The JSCC decoder reverses the encoder’s operations, and its output serves as input to the conditional diffusion decoder, which is built on a U-Net model. Features of different scales, encoded by hierarchical encoders, are combined with time information to guide the denoising process. The conditional diffusion decoder iteratively refines the output until $\bm{\hat{x}}_{0}$ is produced.
@@ III-B Loss Function
For an entropy model-based rate-adaptive JSCC system [6], the loss function is typically defined as:
@@ III-C Training Strategy
To ensure training stability and enhance overall performance, we propose a multi-stage training strategy. Initially, we train each module individually to reduce complexity and facilitate convergence. Once all modules have been trained separately, we fine-tune the entire model. The complete multi-stage training process is detailed in Algorithm 1.
@@
3 First Stage: Train Compression Modules.
@@ 7       Calculate the loss function:
8       $\mathcal{L}=(1-\eta)\mathbb{E}_{\bm{x}_{0}}||\bm{x}_{0}-\bm{\bar{x}}_{0}||^{2}+\eta\mathbb{E}_{\bm{x}_{0}}[d_{p}(\bm{x}_{0},\bm{\bar{x}}_{0})]+\lambda{\mathbb{E}_{\bm{x}_{0}}[-\log p(\bm{\tilde{z}}|\bm{\tilde{y}})-\log p(\bm{\tilde{y}})]}.$
@@
13 Second Stage: Train Transmission Modules.
@@
22 Third Stage: Finetune the whole model.
23 Load the parameters trained in the previous stages.
@@ IV Simulation Results
(a) The performance on Kodak dataset with average $\text{CBR}=1/48$.
(b) The performance on Kodak dataset with average $\text{CBR}=1/24$.
(c) The performance on MS-COCO dataset with average $\text{CBR}=1/48$.
Figure 3: The performance of the proposed schemes versus SNR.
@@ IV-A1 Basic Settings
We evaluate our CDM-JSCC on the Kodak dataset ($24$ images, each with a resolution of $512\times 768$) and MS-COCO dataset (we select $2,695$ test images larger than $512\times 512$). Our CDM-JSCC trained with $\eta=0.5$, primarily optimized for high perceptual quality, is referred to as CDM-JSCC-P. In contrast, the model trained with a lower $\eta=0.1$, aimed at minimizing MSE distortion, is denoted as CDM-JSCC-D. All experiments are conducted using Pytorch. For the training process, we use the Adam optimizer for stochastic gradient descent, starting with a learning rate of $1\times 10^{-4}$ and planning to reduce it after several epochs. The training dataset comprises $50,000$ randomly sampled images from the ImageNet dataset, which are randomly cropped to $256\times 256$. The batch size is set to $4$, with the epochs for each training stage set to $90,45$, and $10$, respectively.
@@ IV-A2 Benchmarks
  • “CDDM”: This scheme utilizes diffusion models to denoise the channel output [11].
  • “DIFFSC”: This method applies diffusion models to enhance images initially reconstructed by DeepJSCC [14].
@@ IV-A3 Considered Metrics
We evaluate the performance of our method along with benchmarks using typical metric PSNR as well as perceptual metrics LPIPS[19], and Fréchet Inception Distance (FID). PSNR calculates pixel-wise MSE distortion. LPIPS measures the $l_{2}$ distance between two latent embeddings extracted by a pre-trained network [19]. FID calculates the distribution of source images and reconstructed images in the feature space using the Fréchet distance [20].
@@ IV-B Performance Comparison
We evaluate the performance of our proposed model on the AWGN channel, a representative and widely adopted choice in this research community. In all subsequent experiments, the model is trained on a single SNR value and tested on the same SNR.
Fig. 3(a) compares CDM-JSCC with benchmarks on the Kodak dataset using an average $\text{CBR}=1/48$ across various SNRs, while Fig. 3(b) presents the results at an average CBR of $1/24$. In addition, Fig. 3(c) shows the results on the MS-COCO dataset at an average CBR of $1/48$. All results demonstrate significant advantages of our proposed CDM-JSCC over the benchmarks, with robust performance across various datasets, CBRs, and SNRs. Notably, our CDM-JSCC-P outperforms all the benchmarks across all perceptual metrics, including LPIPS and FID. This enhancement is attributed to the proposed loss function and the powerful generative diffusion models. However, it is important to note that our CDM-JSCC-P shows performance degradation in the PSNR metric. This is because generative model-based CDM-JSCC-P tends to generate realistic and clear images and may overlook pixel-wise fidelity sometimes. Despite this, our CDM-JSCC-P still outperforms GAN-JSCC in the PSNR metric. Meanwhile, our CDM-JSCC-D is also competitive. Setting $\eta=0.1$ leads to a larger MSE distortion term compared to $\eta=0.5$, thereby achieving better pixel-level fidelity measured by PSNR. As shown in Fig. 3, CDM-JSCC-D not only surpasses DeepJSCC and DIFFSC in PSNR but also offers significantly improved perceptual quality over most benchmarks. Besides, the decoding time of CDM-JSCC is approximately $0.76$s per Kodak image, showing advantages over BPG+LDPC of $4.6$s and CDDM of $1.73$s.